# FreeDrag: Feature Dragging for Reliable Point-based Image Editing

### v3

@@ Abstract
To serve the intricate and varied demands of image editing, precise and flexible manipulation in image content is indispensable. Recently, Drag-based editing methods have gained impressive performance. However, these methods predominantly center on point dragging, resulting in two noteworthy drawbacks, namely “miss tracking”, where difficulties arise in accurately tracking the predetermined handle points, and “ambiguous tracking”, where tracked points are potentially positioned in wrong regions that closely resemble the handle points. To address the above issues, we propose FreeDrag, a feature dragging methodology designed to free the burden on point tracking. The FreeDrag incorporates two key designs, i.e., template feature via adaptive updating and line search with backtracking, the former improves the stability against drastic content change by elaborately controls feature updating scale after each dragging, while the latter alleviates the misguidance from similar points by actively restricting the search area in a line. These two technologies together contribute to a more stable semantic dragging with higher efficiency. Comprehensive experimental results substantiate that our approach significantly outperforms pre-existing methodologies, offering reliable point-based editing even in various complex scenarios.
@@ 1 Introduction
The domain of image editing utilizing generative models has gained substantial attention and witnessed remarkable advancements in recent years [38, 36, 31, 24, 14, 10]. In order to effectively address the intricate and diverse demands of image editing in real-world applications, it becomes imperative to achieve precise and flexible manipulation of image content. Consequently, researchers have proposed two primary categories of methodologies in this domain: (1) harnessing prior 3D models [8, 12, 46] or manual annotations [2, 17, 26, 34, 42] to enhance control over generative models, and (2) employing textual guidance for conditional generative models [37, 39, 41]. Nevertheless, the former category of methodologies often encounter challenges in generalizing to novel assets, while the latter category exhibits limitations in terms of precision and flexibility when it comes to spatial attribute editing.
@@
To overcome these aforementioned limitations, a recent pioneering study, known as DragGAN [33], has emerged as a remarkable contribution in the realm of precise image editing. This work has garnered significant attention, primarily due to its interactive point-based editing capability, termed “drag” editing, which enables users to exert precise control over the editing process by specifying pairs of handle and target points on the given image. The DragGAN framework introduces a two-step iterative process: (i) a motion supervision step, which directs the handle points to migrate towards their corresponding target positions, and (ii) a point tracking step, which consistently tracks the relocated handle points’ positions. In each iteration, the points derived from the current iteration necessitate supervision from points of the last iteration and are subsequently tracked for the next iteration. We categorize this type of method, exemplified by DragGAN and its variants [43], as point dragging solution.
@@
Notwithstanding the praiseworthy achievements exhibited by point dragging solution, there exist several issues. One issue is miss tracking, whereby point dragging encounter difficulty in effectively tracking the desired handle points. This issue arises particularly in highly curved regions with a large perceptual path length, as observed in latent space [21]. In such cases, the optimized image undergoes drastic changes, leading to handle points in subsequent iterations being positioned outside the intended search region. Additionally, in certain scenarios, miss tracking leads to the disappearance of handle points, as shown in Figure 2. It is important to note that during miss tracking, the cumulative error in the motion supervision step increases progressively as iterations proceed, owing to the misalignment of tracked features. Another issue that arises is ambiguous tracking, where the tracked points are situated within other regions that bear resemblance to the handle points. This predicament emerges when there are areas in the image that possess similar features to the intended handle points, leading to ambiguity in the tracking process. (see Figure 3). This issue introduces a potential challenge as it can misguide the motion supervision process in subsequent iterations, leading to inaccurate or misleading directions.
@@
To remedy the aforementioned issues, we propose FreeDrag, a feature dragging solution for interactive point-based image editing. To address the miss tracking issue, we introduce a template feature which is maintained for each handle point to supervise the movements during the iterative process. This template feature is implemented as an exponential moving average feature that dynamically adjusts its weights based on the errors encountered in each iteration. Even when miss tracking occurs in a specific iteration, the maintained template feature remains intact, preventing the optimized image from undergoing drastic changes. To handle the ambiguous tracking issue, we propose line search with backtracking. Line search restricts the movements along a specific line connecting the original handle point and the corresponding target point. This constraint effectively reduces the presence of ambiguous points and minimizes the potential misguidance of the movement direction in subsequent iterations. Moreover, backtracking mechanism enables prompt adjustment for motion plan by effectively discriminating abnormal motion, thereby enhancing the reliability of total movement process. In light of the fact that the points in each iteration undergo supervision from template features and do not necessitate exacting tracking precision, we classify our approach as a feature dragging solution. To summarize, our key contributions are as follows:
  • We have observed that previous drag-based editing approaches adopt a point dragging solution, resulting in miss tracking and ambiguous tracking.
@@
  • We propose FreeDrag, a feature dragging solution for reliable point-based image editing that incorporates adaptive template features and line search with backtracking.
@@
  • We propose FreeDragBench, a new evaluation dataset with 2251 handmade dragging instructions that toiled for GAN-based dragging editing, equipped with a new metric, which measures the editing accuracy of a pair symmetrical dragging instructions.
  • Extensive experiments demonstrate the superiority and stability of FreeDrag in point-based image editing, marking a significant advancement in the field of flexible and precise image editing.
@@ 2.2 Diffusion Models
The emerging diffusion models [15, 44], which conducts gradual denoising procedures from Gaussian noises to nature images, have recently sparked a strong wave of more potent image synthesis. Based on its promising generation capability, a series of versatile methods [25, 3, 51, 20, 47] are developed to exceed the performance peaks of various generation tasks. Typically, Rombach et al. propose Latent Diffusion Model (LDM)[40], which employs a pre-trained auto-encoder for perceptual compression and then performs high-quality sample in latent space, bringing a substantial advancement in high-resolution image synthesis.
@@ 2.3 Point-based Image Editing
Given an image, interactive image editing aims to modify certain image content in response to specific user input, such as text instructions [29, 28, 4, 53], region mask [27], and reference images [49, 5]. The uniqueness of point-based image editing lies on that the user input is a set of point coordinates, and the generative models are expected to achieve precise image content manipulation to match the intent of users. For instance, Endo [10] devises a latent transformer architecture to learn the mapping between two latent codes in StyleGAN. However, this framework necessitates the aid of a pre-trained optical flow network and demands training procedure toiled for each models, which limits its practicability. Later, DragGAN [33] garners considerable attention with remarkable performance, which performs a cycle of point tracking and motion supervision in the feature map to persistently force handle point move to target point. This simple framework achieves impressive performance and attracts subsequent works [43, 32] for better combination with the popular diffusion models. However, the framework of DragGAN is insecure and shoddy edited results may appear. Therefore, how to harness the unstability in dragging process for a more reliable point-based image editing remains to be further explored.
@@ 3 Motivation
Given a set of $n$ handle points $\left\{p_{1},p_{2},p_{3}...,p_{n}\right\}$ and a corresponding set of $n$ target points $\left\{t_{1},t_{2},t_{3}...,t_{n}\right\}$, the objective of point-based dragging is to displace $p_{i}$ to its respective $t_{i}$. Illustrated in Fig. 4, the widely-adopted DragGAN [33] accomplishes this objective through two sequential steps in each motion: (i) Motion Supervision, wherein the current handle point is consistently directed towards its target point by leveraging the feature of itself. (ii) Point Tracking, involving the search for the handle point in the proximity of the handle point from the last motion. Denoting the initial feature map as $F_{0}$, the tracked handle point $p_{i}^{k}$ for the $k$-th motion possesses the most similar feature to $F_{0}(p_{i}^{0})$ in the 2D tracking area centered at $p_{i}^{k-1}$.
@@
While the point dragging pipeline depicted in Fig. 4 presents a promising solution for point-based image editing, it is noted that it frequently encounters challenges, including handle point loss, imprecise editing, and distorted image generation in certain scenarios. These issues are attributed to the intrinsic instability of point dragging, encompassing miss tracking and ambiguous tracking.
(i) Miss Tracking: This occurs in situations where point dragging encounters difficulty in effectively tracking the designated handle points. Given the presence of highly curved regions with substantial perceptual path lengths, as discerned in latent space [21], the optimized image undergoes significant alterations following motion supervision. Consequently, the handle point $p_{i}^{k+1}$ deviates outside the intended search region of $p_{i}^{k}$, as shown in Figure 2, leading to miss tracking in the point tracking step. In specific scenarios, $p_{i}^{k+1}$ may completely vanish from the entire feature map, exemplified by the disappeared glasses in Figure 2. It is imperative to underscore that during miss tracking, the cumulative error in the motion supervision step progressively amplifies with iterations due to the misalignment of tracked features. (ii) Ambiguous Tracking: This occurs when the tracked points are positioned within other regions that bear resemblance to the handle points. This challenge arises when there are areas in the image exhibiting features similar to the intended handle points, such as the blue boundary lines and horse’s hooves in Figure 3, which may misdirect the motion supervision process in subsequent iterations, resulting in inaccurate or misleading directional adjustments.
@@ 4 Methodology
In light of the instability associated with point dragging, which heavily depends on accurate point tracking in each step, we introduce a feature dragging approach termed FreeDrag, as illustrated in Fig. 5(a). Here, $h_{i}^{k}$ represents the target position in the $k$-th drag, and $F_{r}(h_{i}^{k})$ signifies the feature aggregate centered at $h_{i}^{k}$ with a radius $r$ in the feature map $F$, i.e.,
$\vspace{-2mm}F_{r}({h_{i}^{k}})=\sum_{q_{i}\in\Omega(h_{i}^{k},r)}F(q_{i}),$ (1)
Here, $\Omega(h_{i}^{k},r)$ denotes the square patch centered at $h_{i}^{k}$ with a side length of $2r$. In the $k$-th drag, we promote $h_{i}^{k}$ to be the carrier of $T_{i}^{k}$ by compelling the feature aggregate $F_{r}(h_{i}^{k})$ to closely align with the template feature $T_{i}^{k}$ (as depicted by the red line in Fig. 5(a)), which can be formulated as:
${\mathcal{L}_{drag}}=\sum_{i=1}^{n}{\left\|F_{r}(h_{i}^{k})-T_{i}^{k}\right\|_{1}}.$ (2)
@@ 4.1 Template Features via Adaptive Updating
$T_{i}^{k+1}=\lambda_{i}^{k}\cdot{F_{r}}(h_{i}^{k})+(1-\lambda_{i}^{k})\cdot T_{i}^{k}.$ (3)
@@
$\lambda_{i}^{k}={(1+exp(\alpha\cdot(L_{(i,k)}^{en}-\beta)))^{-1}},$ (4)
@@
$0.5={(1+exp(\alpha\cdot(0.2\cdot l-\beta)))^{-1}},$ (5)
$0.1={(1+exp(\alpha\cdot(0.8\cdot l-\beta)))^{-1}}.$ (6)
@@ 4.2 Line Search with Backtracking
$=\operatorname*{arg\,min}_{q_{i}\in\pi(h_{i}^{k},t_{i},d)}{\left\|{||{F_{r}}(q_{i})-T_{i}^{k+1}|{|_{1}}}\right.{\left.{-l}\right\|_{1}}}$ (8)
@@
Additionally, as depicted in Fig. 5(b), during the joint optimization of multiple point dragging, the motion direction of a specific point may be influenced by the overall trend. This can result in the handle point deviating from the target point in certain steps. For instance, in comparison to $p_{2}^{0}$, the handle point $p_{2}^{1}$ is farther away from $h_{2}^{1}$. To address this issue, we integrate a backtracking mechanism to identify such abnormal movements, facilitating prompt adjustments for the subsequent dragging plan. Concretely, backtracking is implemented by introducing two additional options for the dragging plan: frozen and fallback the point, which can be expressed as:
$h_{i}^{k+1}=\left\{\begin{array}[]{l}{S}(h_{i}^{k},t_{i},T_{i}^{k+1},d,l),\quad if\ L_{(i,k)}^{en}\leq 0.5\cdot{l}\\ \ \ \ \ \ \ h_{i}^{k},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \quad\ elif\ \ \ L_{(i,k)}^{en}\leq L_{(i,k)}^{in}\\ S(h_{i}^{k}-d\cdot\frac{{{t_{i}}-h_{i}^{k}}}{{{{\left\|{{t_{i}}-h_{i}^{k}}\right\|}_{2}}}},t_{i},T_{i}^{k+1},2d,0),else\end{array}\right.$ (9)
@@ 4.3 Termination Signal
For each feature dragging towards $h_{i}^{k}$ , the maximum optimization step of each feature dragging is set as 5. To enhance efficiency, we pause the optimization process if $L_{(i,k)}^{en}$ already falls below $0.5\cdot l$. The final termination signal is obtained by determining that if the remaining distance $||h_{i}^{k}-t_{i}||_{2}\leq 2$.

### v4

@@ Abstract
To serve the intricate and varied demands of image editing, precise and flexible manipulation in image content is indispensable. Recently, Drag-based editing methods have gained impressive performance. However, these methods predominantly center on point dragging, resulting in two noteworthy drawbacks, namely “miss tracking”, where difficulties arise in accurately tracking the predetermined handle points, and “ambiguous tracking”, where tracked points are potentially positioned in wrong regions that closely resemble the handle points. To address the above issues, we propose FreeDrag, a feature dragging methodology designed to free the burden on point tracking. The FreeDrag incorporates two key designs, i.e., template feature via adaptive updating and line search with backtracking, the former improves the stability against drastic content change by elaborately controlling the feature updating scale after each dragging, while the latter alleviates the misguidance from similar points by actively restricting the search area in a line. These two technologies together contribute to a more stable semantic dragging with higher efficiency. Comprehensive experimental results substantiate that our approach significantly outperforms pre-existing methodologies, offering reliable point-based editing even in various complex scenarios.
@@
11footnotetext:  indicates equal contributions.22footnotetext:  Corresponding authors.
@@ 1 Introduction
The domain of image editing utilizing generative models has gained substantial attention and witnessed remarkable advancements in recent years [38, 36, 31, 24, 14, 10]. In order to effectively address the intricate and diverse demands of image editing in real-world applications, it becomes imperative to achieve precise and flexible manipulation of image content. Consequently, researchers have proposed two primary categories of methodologies in this domain: (1) harnessing prior 3D models [8, 12, 46] or manual annotations [2, 17, 26, 34, 42] to enhance control over generative models, and (2) employing textual guidance for conditional generative models [37, 39, 41]. Nevertheless, the former category of methodologies often encounters challenges in generalizing to novel assets, while the latter category exhibits limitations in terms of precision and flexibility when it comes to spatial attribute editing.
@@
To tackle these aforementioned limitations, a recent pioneering study, known as DragGAN [33], has emerged as a remarkable contribution in the realm of precise image editing. This work has garnered significant attention, primarily due to its interactive point-based editing capability, termed “drag” editing, which enables users to exert precise control over the editing process by specifying pairs of handle and target points on the given image. The DragGAN framework introduces a two-step iterative process: (i) a motion supervision step, which directs the handle points to migrate towards their corresponding target positions, and (ii) a point tracking step, which consistently tracks the relocated handle points’ positions. In each iteration, the points derived from the current iteration necessitate supervision from points of the last iteration and are subsequently tracked for the next iteration. We categorize this type of method, exemplified by DragGAN and its variant [43], as point dragging solutions.
@@
Notwithstanding the praiseworthy achievements exhibited by point dragging solution, there exist several issues. One issue is miss tracking, whereby point dragging encounters difficulty in effectively tracking the desired handle points. This issue arises particularly in highly curved regions with a large perceptual path length, as observed in latent space [21]. In such cases, the optimized image undergoes drastic changes, leading to handle points in subsequent iterations being positioned outside the intended search region. Additionally, in certain scenarios, miss tracking leads to the disappearance of handle points, as shown in Figure 2. It is important to note that during miss tracking, the cumulative error in the motion supervision step increases progressively as iterations proceed, owing to the misalignment of tracked features. Another issue that arises is ambiguous tracking, where the tracked points are situated within other regions that bear resemblance to the handle points. This predicament emerges when there are areas in the image that possess similar features to the intended handle points, leading to ambiguity in the tracking process. (see Figure 3). This issue introduces a potential challenge as it can misguide the motion supervision process in subsequent iterations, leading to inaccurate or misleading directions.
@@
To remedy the aforementioned issues, we propose FreeDrag, a feature dragging solution for interactive point-based image editing. To address the miss tracking issue, we introduce a template feature that is maintained for each handle point to supervise the movements during the iterative process. This template feature is implemented as an exponential moving average feature that dynamically adjusts its weights based on the errors encountered in each iteration. Even when miss tracking occurs in a specific iteration, the maintained template feature remains intact, preventing the optimized image from undergoing drastic changes. To handle the ambiguous tracking issue, we propose the line search with backtracking. Line search restricts the movements along a specific line connecting the original handle point and the corresponding target point. This constraint effectively reduces the presence of ambiguous points and minimizes the potential misguidance of the movement direction in subsequent iterations. Moreover, the backtracking mechanism enables prompt adjustment for motion plan by effectively discriminating abnormal motion, thereby enhancing the reliability of the total movement process. In light of the fact that the points in each iteration undergo supervision from template features and do not necessitate exacting tracking precision, we classify our approach as a feature dragging solution. To summarize, our key contributions are as follows:
@@
  • We propose FreeDrag, a feature dragging solution for reliable point-based image editing that incorporates adaptive template features and line search with backtracking, marking a significant advancement in the field of flexible and precise image editing.
@@
  • We propose FreeDragBench, a new evaluation dataset with 2251 handmade dragging instructions that are tailored for GAN-based dragging editing, equipped with a new metric, which measures the editing accuracy of a pair of symmetrical dragging instructions.
@@ 2.2 Diffusion Models
The emerging diffusion models [15, 44], which conduct gradual denoising procedures from Gaussian noises to natural images, have recently sparked a strong wave of more potent image synthesis. Based on its promising generation capability, a series of versatile methods [25, 3, 51, 20, 47] are developed to exceed the performance peaks of various generation tasks. Typically, Rombach et al. propose the Latent Diffusion Model (LDM)[40], which employs a pre-trained auto-encoder for perceptual compression and then performs high-quality sample in latent space, bringing a substantial advancement in high-resolution image synthesis.
@@ 2.3 Point-based Image Editing
Given an image, interactive image editing aims to modify certain image content in response to specific user input, such as text instructions [29, 28, 4, 53], region mask [27], and reference images [49, 5]. The uniqueness of point-based image editing lies in that the user input is a set of point coordinates, and the generative models are expected to achieve precise image content manipulation to match the intent of users. For instance, Endo [10] devises a latent transformer architecture to learn the mapping between two latent codes in StyleGAN. However, this framework necessitates the aid of a pre-trained optical flow network and demands a training procedure tailored for each model, which limits its practicability. Later, DragGAN [33] garners considerable attention with remarkable performance, which performs a cycle of point tracking and motion supervision in the feature map to persistently force the handle point to move to the target point. This simple framework achieves impressive performance and attracts subsequent works [43, 32] for better combination with the popular diffusion models.
Generally, the GAN-based dragging approaches achieve superior dragging compared to diffusion-based approaches but exhibit inferior real image inversion. The GAN-based approaches benefit from the attribute disentanglement of StyleGAN, enhancing dragging capability. However, its generative quality and real image inversion ability are comparatively limited. In contrast, diffusion models achieve higher generative quality and superior real image inversion. Nevertheless, it encounters challenges in balancing point manipulation and appearance preservation due to the intertwined feature map, and demands more processing time.
@@ 3 Motivation
Given a set of $n$ handle points $\left\{p_{1},p_{2},p_{3}...,p_{n}\right\}$ and a corresponding set of $n$ target points $\left\{t_{1},t_{2},t_{3}...,t_{n}\right\}$, the objective of point-based dragging is to displace $p_{i}$ to its respective $t_{i}$. Illustrated in Fig. 4, the widely adopted DragGAN [33] accomplishes this objective through two sequential steps in each motion: (i) Motion Supervision, wherein the current handle point is consistently directed towards its target point by leveraging the feature of itself. (ii) Point Tracking, involving the search for the handle point in the proximity of the handle point from the last motion. Denoting the initial feature map as $F_{0}$, the tracked handle point $p_{i}^{k}$ for the $k$-th motion possesses the most similar feature to $F_{0}(p_{i}^{0})$ in the 2D tracking area centered at $p_{i}^{k-1}$.
@@
While the point dragging pipeline depicted in Fig. 4 presents a promising solution for point-based image editing, it is noted that it frequently encounters challenges, including handle point loss, imprecise editing, and distorted image generation in certain scenarios. These issues are attributed to the intrinsic instability of point dragging, encompassing miss tracking and ambiguous tracking. (i) Miss Tracking: This occurs in situations where point dragging encounters difficulty in effectively tracking the designated handle points. Given the presence of highly curved regions with substantial perceptual path lengths, as discerned in latent space [21], the optimized image undergoes significant alterations following motion supervision. Consequently, the handle point $p_{i}^{k+1}$ deviates outside the intended search region of $p_{i}^{k}$, as shown in Figure 2, leading to miss tracking in the point tracking step. In specific scenarios, $p_{i}^{k+1}$ may completely vanish from the entire feature map, exemplified by the disappeared glasses in Figure 2. It is imperative to underscore that during miss tracking, the cumulative error in the motion supervision step progressively amplifies with iterations due to the misalignment of tracked features. (ii) Ambiguous Tracking: This occurs when the tracked points are positioned within other regions that bear resemblance to the handle points. This challenge arises when there are areas in the image exhibiting features similar to the intended handle points, such as the blue boundary lines and horse’s hooves in Figure 3, which may misdirect the motion supervision process in subsequent iterations, resulting in inaccurate or misleading directional adjustments.
@@ 4 Methodology
In light of the instability associated with point dragging, which heavily depends on accurate point tracking in each step, we introduce a feature dragging approach termed FreeDrag, as illustrated in Fig. 5(a). Here, $h_{i}^{k}$ represents the target position in the $k$-th drag, and $F_{r}(h_{i}^{k})$ signifies the feature aggregate centered at $h_{i}^{k}$ with a radius $r$ in the feature map $F$, which can be expressed as:
$F_{r}({h_{i}^{k}})=\sum_{q_{i}\in\Omega(h_{i}^{k},r)}F(q_{i}).\vspace{-3mm}$ (1)
Here, $\Omega(h_{i}^{k},r)$ denotes the square patch centered at $h_{i}^{k}$ with a side length of $2r$. In the $k$-th drag, we promote $h_{i}^{k}$ to be the carrier of $T_{i}^{k}$ by compelling the feature aggregate $F_{r}(h_{i}^{k})$ to closely align with the template feature $T_{i}^{k}$ (as depicted by the red line in Fig. 5(a)), i.e.,
${\mathcal{L}_{drag}}=\sum_{i=1}^{n}{\left\|F_{r}(h_{i}^{k})-T_{i}^{k}\right\|_{1}}.\vspace{-2mm}$ (2)
@@ 4.1 Template Features via Adaptive Updating
$T_{i}^{k+1}=\lambda_{i}^{k}\cdot{F_{r}}(h_{i}^{k})+(1-\lambda_{i}^{k})\cdot T_{i}^{k}.\vspace{-2mm}$ (3)
@@
$\lambda_{i}^{k}={(1+exp(\alpha\cdot(L_{(i,k)}^{en}-\beta)))^{-1}},\vspace{-2mm}$ (4)
@@
$0.5={(1+exp(\alpha\cdot(0.2\cdot l-\beta)))^{-1}},\vspace{-1mm}$ (5)
$0.1={(1+exp(\alpha\cdot(0.8\cdot l-\beta)))^{-1}}.\vspace{-2mm}$ (6)
@@ 4.2 Line Search with Backtracking
$=\mathop{\arg\min}\limits_{{q_{i}\in\pi(h_{i}^{k},t_{i},d)}}\left\lVert\lVert F_{r}(q_{i})-T_{i}^{k+1}\rVert_{1}-l\right\rVert_{1},$ (8)
@@
Additionally, as depicted in Fig. 5(b), during the joint optimization of multiple points dragging, the motion direction of a specific point may be influenced by the overall trend. This can result in the handle point deviating from the target point in certain steps. For instance, in comparison to $p_{2}^{0}$, the handle point $p_{2}^{1}$ is farther away from $h_{2}^{1}$. To address this issue, we integrate a backtracking mechanism to identify such abnormal movements, facilitating prompt adjustments for the subsequent dragging plan. Concretely, backtracking is implemented by introducing two additional options for the dragging plan: frozen and fallback the point, which can be expressed as:
$h_{i}^{k+1}\!=\!\!\left\{\!\!\begin{array}[]{l}{S}(h_{i}^{k},t_{i},T_{i}^{k+1},d,l),\quad if\ L_{(i,k)}^{en}\leq 0.5\cdot{l}\\ \ \ \ \ \ \ h_{i}^{k},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \quad\ elif\ \ \ L_{(i,k)}^{en}\leq L_{(i,k)}^{in}\\ S(h_{i}^{k}\!-\!d\cdot\frac{{{t_{i}}-h_{i}^{k}}}{{{{\left\|{{t_{i}}-h_{i}^{k}}\right\|}_{2}}}},t_{i},T_{i}^{k+1},2d,0),else\end{array}\right.\vspace{-2mm}$ (9)
@@ 4.3 Termination Signal
For each feature dragging towards $h_{i}^{k}$ , the maximum optimization step of each feature dragging is set as 5. To enhance efficiency, we pause the optimization process if $L_{(i,k)}^{en}$ already falls below $0.5\cdot l$. The final termination signal is obtained by determining if the remaining distance $||h_{i}^{k}-t_{i}||_{2}\leq 2$.